A bank teller has 54 $5 and $20 bills in her drawer. The value of the bills is $780. How many $5 bills are there?
Answer by swincher4391(726)

f + t = 54
5f + 20t = 780
================
20f +20t = 1080
5f + 20t = 780
15f = 300
f = 20
============
You have 20 fives.